# From Cavitation to Astrophysics: Explicit Solution of the Spherical Collapse Equation

### 12 Jan 2024 v2

@@ Abstract
Differential equations of the form $\ddot{R}=-kR^{\gamma}$, with a positive constant $k$ and real parameter $\gamma$, are fundamental in describing phenomena such as the spherical gravitational collapse ($\gamma=-2$), the implosion of cavitation bubbles ($\gamma=-4$) and the orbital decay in binary black holes ($\gamma=-7$). While explicit elemental solutions exist for select integer values of $\gamma$, more comprehensive solutions encompassing larger subsets of $\gamma$ have been independently developed in hydrostatics (see Lane-Emden equation) and hydrodynamics (see Rayleigh-Plesset equation). This paper introduces a general explicit solution for all real $\gamma$, employing the quantile function of the beta distribution, readily available in most modern programming languages. This solution bridges between distinct fields and reveals insights, such as a critical branch point at $\gamma=-1$, thereby enhancing our understanding of these pervasive differential equations.
keywords:
gravitation, hydrodynamics, black hole physics, cosmology: theory
1 Introduction
@@
with fixed $k\in\mathbb{R}_{+}$ and $\gamma\in\mathbb{R}$. Dots denote derivatives with respect to time $T$. Hereafter, I call Equation (1) the ‘spherical collapse’ equation as it is often used in this context.
I will restrict the discussion to the case, where $R$ exhibits a finite maximum $R_{0}$ at some finite time—in many applications this is equivalent to stating that the system is ‘bound’ with negative total energy. Equation (1) being an autonomous ODE, we can always choose the origin of time to coincide with $R=R_{0}$, in which case the initial conditions read
@@
Table 1 lists various examples of physical systems governed by Equation (1). One of the most famous cases in astrophysics is the gravitational collapse of a uniform collisionless sphere ($\gamma=-2$), the so-called top-hat spherical collapse model (Gunn & Gott, 1972). A simple parametric solution, $\{T(\theta),R(\theta)\}$ (Lin et al., 1965), has become the default textbook solution. Its transcendental nature implies that algebraic solutions of Equation (1) do not generally exist, yet the search for explicit non-algebraic solutions and fast approximations continues, e.g., Slepian & Philcox (2023), who applied Fourier expansions to complex contour integration.
Barely known to astrophysicists, an analogous collapse equation (but with $\gamma=-4$) has long been studied in hydrodynamics. This equation, dating back to Stokes (1847) and Rayleigh (1917), describes the collapse of an empty spherical cavity in an ideal incompressible liquid. Approximate solutions $R(T)$ derived by Obreschkow et al. (2012) have led Amore & Fernández (2013) to develop an infinite series that rapidly converges towards the exact solution. Shortly after, Kudryashov & Sinelshchikov (2014, 2015) presented the first closed-form solution $T(R)$ for cavities in $N\geq 3$ dimensions, in terms of hypergeometric functions. These can, in principle, be inverted to explicit solutions $R(T)$.
Equation (1) is also encountered in static systems. For example, it is analogous to the one-dimensional Lane-Emden equation (Lane, 1870; Emden, 1907), describing the density profile of a self-gravitating polytropic gas in a thin tube in hydrostatic equilibrium. A solution for $\gamma\!>\!-1$ was presented by Harrison & Lake (1972) in the context of galactic discs.
Building on the special solutions found independently in different fields, this paper presents a general explicit solution of Equation (1) for any real $\gamma$. This solution constitutes a mathematical unification of a wide range of unrelated topics in physics and astrophysics. Section 2 overviews basic reformulations, useful for the derivation of the general solution in Section 3 and its discussion in Section 4. Section 5 concludes with a brief synthesis and outlook.
@@
[…]
Generalisation to a spherical bubble in $N\geq 3$ dimensions (e.g., Klotz, 2013)
@@
Two-body collision Two point masses in gravitational free-fall towards each other, on a straight line without angular momentum
@@
Free-fall with constant acceleration Straight fall in a constant acceleration field, such as a vertical drag-free drop of an object on Earth
@@
Table 1: Examples of systems governed by Equations (1) and (2). Moving objects described by $R(T)$ are shown in blue, whereas static objects described by this equation are shown in red. $T_{0}$ is the characteristic physical scale (normally a characteristic time), such that Equation (1) reduces to the dimensionless Equation (5) upon normalisation via Equation (3). The collapse time is $T_{c}$, satisfying $R(T_{c})=0$, is given by $T_{0}\tau$. Where $T_{0}$ depends on $R_{0}$, the latter is the maximum, initial value of $R$. In the case of the polytrope (bottom row), $R$ and $T$ have different meanings: $R:=\rho^{1/n}$ is a measure of the gas density and $T$ is the distance from the centre of mass. For most examples, analogous cases with other forces, such as electric forces instead of gravitational ones, can be found.
2 Setting the stage
Before delving into the solutions of Equation (1), it is worth stating a few straightforward, but pivotal remarks on equivalent reformulations. These will help the derivation (Section 3) and discussion (Section 4) of the solutions.
2.1 Dimensionless from
The first remark is that it is often more convenient to work in dimensionless position and time coordinates, defined as
@@ 2.1 Dimensionless from
The constant $k$ has conveniently disappeared in the dimensionless form of Equation (5), reducing the family of ODEs to a single control parameter $\gamma$. The natural disappearance of $k$ brings to light the inherently scale-free nature of Equation (1) implied by its power law structure. This scale-invariance is the deeper reason for Equation (1) to be applicable from microscopic to astrophysical scales.
As a simple example, let us consider the gravitational collapse of a uniform pressure-free sphere, initially at rest with radius $R(0)=R_{0}$. The Newtonian equation of motion,
@@
is identical to Equation (1) with constants $\gamma=-2$ and $k=GM$. Hence, the natural time constant is $T_{0}=\sqrt{R_{0}^{3}/(GM)}$. Upon expressing $R$ and $T$ in their natural units $R_{0}$ and $T_{0}$, Equation (7) becomes indeed $\ddot{r}=-r^{-2}$.
For convenience, the remainder of this paper will focus on dimensionless solutions $r(t)$. Corresponding dimensional solutions $R(T)$ are easily obtained using (Equation 3), given the scales $R_{0}$ and $T_{0}$. Explicit expressions for these scales in various physical systems are provided in Table 1.
2.2 Equivalent formulations
@@ 2.2.1 Integral of motion
If differentiated with respect to time and solved for $\ddot{r}$, Equation (8b) becomes Equation (5). Strictly, this is true if and only if $\dot{r}\neq 0$, as the transition involves a division by $\dot{r}$. Since Equations (8b) are first-order differential equations, only one boundary condition is required, such as $r(0)=1$.
2.2.2 Hybrid differential equation
If $\gamma\neq-1$, a hybrid differential equation, mixing first and second derivatives, can be obtained by rewriting $r^{1+\gamma}$ in Equation (8a) as $rr^{\gamma}=-r\ddot{r}$ (Equation 5), hence
@@ 2.2.2 Hybrid differential equation
This is, for example, the standard form of the equation describing the collapse of a spherical cavitation bubble without viscosity and surface tension (Rayleigh, 1917). This equation is normally written in dimensional form,
@@
where $\Delta p$ is the driving pressure and $\rho$ the constant liquid density (see first row in Table 1). If normalising $R$ by $R_{0}=R(0)$ and $T$ by $T_{0}=R_{0}\sqrt{\rho/\Delta p}$, Equation (10) becomes Equation (9) with $\gamma=-4$. It follows that Equation (10) is equivalent to Equation (1) with $\gamma=-4$ and $k=R_{0}^{3}\Delta p/\rho$.
Figure 1: Potential $\phi(r)$, given in Equation (12), for different values of $\gamma$, equally spaced by $\Delta\gamma=0.2$. There are qualitatively distinct families at $\gamma<-1$ (red–yellow) and $\gamma>-1$ (blue–green), as discussed in Section 4.1. They are separated by an isolated critical curve for $\gamma=-1$. Thin black lines highlight the coloured curves with a printed $\gamma$-value.
2.2.3 Lagrangian
@@
Applying the Euler-Lagrange equation, $\tfrac{{\rm d}}{{\rm d}t}(\partial\mathcal{L}/\partial\dot{r})=\partial\mathcal{L}/\partial r$, generates Equation (5), for all $\gamma$.
Selected potentials $\phi(r)$ are shown in Figure 1. The existence of two regimes, separated by the critical value $\gamma=-1$, is the fundamental reason for the qualitatively different behaviour of $r(t)$ in these regimes. These regimes will be discussed in more detail in Section 4.1.
3 General solution
In this section, I derive compact general solutions of Equation (1) with initial conditions of Equation (2), following an approach similar to Kudryashov & Sinelshchikov (2014), but for general real parameters $\gamma$. For convenience, all derivations and solutions are presented for the dimensionless form in Equation (5) with initial conditions stated in Equation (6). Any solution $r(t)$ can be transformed to the corresponding dimensional solution $R(T)$ via the straightforward linear transformations of Equation (3).
3.1 Implicit solution $t(r)$
3.1.1 Cases $\gamma\neq-1$
${\rm d}t^{2}=\frac{1+\gamma}{2(1-r^{1+\gamma})}{\rm d}r^{2}.$ (13)
@@ 3.1.1 Cases $\gamma\neq-1$
$\alpha=\frac{1}{4}+\frac{3-\gamma}{4|1+\gamma|}.$ (19)
We have defined $\alpha$ in this way to make the right-hand integral of Equation (18) identical to the definition of the incomplete beta function, $B(x;\alpha,\beta)=\int_{0}^{x}y^{\alpha-1}(1-y)^{\beta-1}{\rm d}y$. Hence, Equation (18) solves to
$t(r)=\tau-\sqrt{\frac{\eta}{2}}~{}B\left(r^{|1+\gamma|};\alpha,\tfrac{1}{2}\right).$ (20)
The collapse time $\tau$ can be obtained by solving this equation for $\tau$ at the point $t=0$ ($r=1$),
$\tau(\gamma)=\sqrt{\frac{\eta}{2}}~{}B(\alpha,\tfrac{1}{2}).$ (21)
Here $B(\alpha,\beta)$ with only two arguments is the (complete) beta function, i.e., the incomplete $B(x;\alpha,\beta)$ evaluated at $x=1$. Selected explicit values of $\tau$ are listed in Table 2.
Substituting Equation (21) back into Equation (20), the latter can be slightly compacted to
$t(r)=\sqrt{\frac{\eta}{2}}~{}B\left(1-r^{|1+\gamma|};\tfrac{1}{2},\alpha\right),$ (22)
analogous to the solution for $\gamma>-1$ by Harrison & Lake (1972), but now generalised to all $\gamma\neq-1$.
@@
Table 2: List of dimensionless collapse times $\tau$, given by Equations (21) and (24), and dimensionless collapse point velocities $\dot{r}(\tau)$, given by Equation (29). Exact solutions are listed if they can be reduced to an algebraic expression with basic constants.
3.1.2 Special case $\gamma=-1$
What remains is the special case of $\gamma=-1$. Equation (8b) describing this case integrates to
$t(r)=\tau(-1)\,{\rm erf}\Big{(}\sqrt{-\ln r}\,\Big{)}$ (23)
where
$\tau(-1)=\sqrt{\pi/2}$ (24)
is the collapse time, reached for $r\rightarrow 0_{+}$. This collapse time is equal to the limit of Equation (21) for $\gamma\rightarrow-1$. Likewise, Equation (23) is the limit of Equation (22) for $\gamma\rightarrow-1$.
3.2 Explicit solution $r(t)$
In order to invert the implicit solutions $t(r)$, it is convenient to normalise Equation (20) by $\tau$, which yields
$\frac{t}{\tau}=1-I\left(r^{|1+\gamma|};\alpha,\tfrac{1}{2}\right),$ (25)
where $I(x;\alpha,\beta)\equiv B(x;\alpha,\beta)/B(\alpha,\beta)$ is called the regularised incomplete beta function (a convenient step pointed out by Roberto Iacono, priv. com.).
The implicit Equations (25) and (23) can be readily inverted to a compact general explicit solution in terms of well-known special functions,
$r(t)=$ $\exp\left(-{\rm erfi}^{2}\Big{(}\sqrt{2/\pi}~{}t\Big{)}\right)$ if $\gamma\!=\!-1$ (26b)
where ${\rm erfi}(x)$ is the inverse error function and $Q(x;\alpha,\beta)$ is the inverse regularised incomplete beta function. As required, Equation (26b) is the limit of Equation (26a) for $\gamma\rightarrow-1$.
The absolute value $|t|$ in Equation (26a) does not follow from Equation (25), only valid for $t\in[0,\tau]$. It is an ad-hoc extension of $r(t)$ to the domain $t\in[-\tau,\tau]$, exploiting the time-reversal symmetry of Equations (5) and (6). Equation (26b) is already time-symmetric. Hence, the solution of Equation (26b) is valid on the interval $t\in[-\tau,\tau]$.
The practical value of Equation (26b) becomes apparent when realising that the regularised incomplete beta function $I(x;\alpha,\beta)$ is the cumulative density function of the beta distribution, a common distribution function in probability theory, related to, but not to be confused with, the beta function. Hence, the function $Q(x;\alpha,\beta)$ is the quantile function of the beta distribution—thus the letter $Q$. This quantile function is available in all modern computing languages (see Section 3.3 for some examples).
Figure 2 shows $r(t)$ on the interval $t\in[0,\tau(\gamma)]$ for all integer $\gamma$ from $-4$ to $+4$. Key properties of these solutions will be discussed in Section 4.

### 19 Mar 2024 v3

@@ Abstract
Differential equations of the form $\ddot{R}=-kR^{\gamma}$, with a positive constant $k$ and real parameter $\gamma$, are fundamental in describing phenomena such as the spherical gravitational collapse ($\gamma=-2$), the implosion of cavitation bubbles ($\gamma=-4$) and the orbital decay in binary black holes ($\gamma=-7$). While explicit elemental solutions exist for select integer values of $\gamma$, more comprehensive solutions encompassing larger subsets of $\gamma$ have been independently developed in hydrostatics (see Lane-Emden equation) and hydrodynamics (see Rayleigh-Plesset equation). I here present a universal explicit solution for all real $\gamma$, invoking the beta distribution. Although standard numerical ODE solvers can readily evaluate more general second order differential equations, this explicit solution reveals a hidden connection between collapse motions and probability theory that enables further analytical manipulations, it conceptually unifies distinct fields, and it offers insights into symmetry properties, thereby enhancing our understanding of these pervasive differential equations.
I Introduction
@@
with fixed $k\in\mathbb{R}_{+}$ and $\gamma\in\mathbb{R}$. Dots denote derivatives with respect to time $T$. Using a capital $T$, normally reserved for temperature, will help distinguish this time from a dimensionless time $t$ (Section II.1). We may call Equation (1) the ‘spherical collapse equation’ by virtue of its most common applications.
I will restrict the discussion to the case, where $R$ exhibits a finite maximum $R_{0}$—in many physical applications this is equivalent to stating that the system is ‘bound’. Equation (1) being an autonomous ODE, we can always choose the origin of time to coincide with $R=R_{0}$, in which case the initial conditions read
@@
Table 1 lists textbook examples of physical systems governed by Equation (1). Many of them are encountered at undergraduate levels and used to illustrate cases, whose solution $R(T)$ must be obtained via numerical ODE solvers, such as Runge–Kutta solvers. The existence, let alone the form, of an explicit solution is rarely noted, except, of course, in trivial cases like the linear free-fall ($\gamma=0$) and harmonic oscillator ($\gamma=1$).
An iconic example of Equation (1) in astrophysics is the gravitational collapse of a uniform collisionless sphere ($\gamma=-2$), the so-called top-hat spherical collapse model [1]. A simple parametric solution, $\{T(\theta),R(\theta)\}$ [2], has become the default solution in cosmology. Its transcendental nature implies that algebraic solutions of Equation (1) do not generally exist, yet the search for non-algebraic [3] and approximate [4] solutions continues.
Barely known to astrophysicists, an analogous collapse equation (but with $\gamma=-4$) has long been studied in hydrodynamics [5]. This equation, named after Lord Rayleigh [6], describes the collapse of an empty spherical cavity in an ideal incompressible liquid. Recent approximations [7] have led to the development of an infinite series that rapidly converges toward the exact solution [8]. Shortly after, others presented the first closed-form solution for cavities in three [9] and $N\geq 3$ [10] dimensions, in terms of hypergeometric functions.
Equation (1) is also encountered in static systems. For example, it is equivalent to the one-dimensional Lane-Emden equation [11, 12], describing the density profile of a self-gravitating polytropic gas in a thin tube in hydrostatic equilibrium. A classic solution for $\gamma\!>\!-1$ has been found in the context of galactic discs [13].
Building on the special solutions found independently in different fields, this paper presents a compact general explicit solution of Equation (1). While overall straightforward, the path to this solution involves some subtleties that enable the generalisation to all real $\gamma$. This allows us to discuss various seemingly unrelated basic problems in physics and astrophysics in unison. Section II overviews reformulations, useful for the derivation of the general solution in Section III and its discussion in Section IV. Section V concludes with a brief synthesis and outlook.
@@
Generalisation to a spherical bubble in $N\geq 3$ dimensions [e.g., 14]
@@
Two-body collision Two point masses in gravitational free-fall toward each other, on a straight line without angular momentum
@@
Free-fall in uniform field Straight fall in a constant acceleration field, such as a vertical drag-free drop of an object on Earth
@@
Table 1: Textbook examples of systems governed by Equations (1) and (2). Moving objects described by $R(T)$ are shown in blue, whereas static objects described by this equation are shown in red. $T_{0}$ is the characteristic physical scale (normally a characteristic time), such that Equation (1) reduces to the dimensionless Equation (5) upon normalisation via Equation (3). The collapse time is $T_{c}$, satisfying $R(T_{c})=0$, is given by $T_{0}\tau$. Where $T_{0}$ depends on $R_{0}$, the latter is the maximum, initial value of $R$. In the case of the polytrope (bottom row), $R$ and $T$ have different meanings: $R:=\rho^{1/n}$ is a measure of the gas density and $T$ is the distance from the centre of mass. For most examples, analogous cases with other forces, such as electric forces instead of gravitational ones, can be found.
II Setting the stage
Before delving into the solutions of Equation (1), it is worth recalling basic, but pivotal remarks on equivalent reformulations. These will help the derivation (Section III) and discussion (Section IV) of the solutions.
II.1 Dimensionless form
The first remark is that it is convenient to work in dimensionless position and time coordinates, defined as
@@ II.1 Dimensionless form
The constant $k$ has conveniently disappeared in the dimensionless form of Equation (5), reducing the family of ODEs to a single control parameter $\gamma$. The natural disappearance of $k$ brings to light the inherently scale-free nature of Equation (1) implied by its power law structure. This is the deeper reason for Equation (1) to be applicable from microscopic to astrophysical scales.
As a textbook example, let us consider the gravitational collapse of a uniform pressure-free sphere, initially at rest with radius $R(0)=R_{0}$. The Newtonian equation of motion reads
@@
which is Equation (1) with $\gamma=-2$ and $k=GM$. Upon expressing $R$ and $T$ in their natural units $R_{0}$ and $T_{0}=\sqrt{R_{0}^{3}/(GM)}$, Equation (7) becomes indeed $\ddot{r}=-r^{-2}$.
II.2 Equivalent formulations
@@ II.2.1 Integral of motion
If differentiated with respect to $t$ and solved for $\ddot{r}$, Equation (8b) becomes Equation (5), except if $\dot{r}=0$. Since Equations (8b) are first-order differential equations, only one boundary condition is required, such as $r(0)=1$.
II.2.2 Hybrid differential equation
If $\gamma\neq-1$, a hybrid ODE, mixing first and second derivatives, can be obtained by rewriting $r^{1+\gamma}$ in Equation (8a) as $rr^{\gamma}=-r\ddot{r}$ (Equation 5), hence
@@ II.2.2 Hybrid differential equation
This is, for example, the standard form of the equation describing the collapse of a spherical cavitation bubble without viscosity and surface tension [6], normally written in dimensional form,
@@
where $\Delta p$ is the driving pressure and $\rho$ the constant liquid density (see first row in Table 1). Expressed in natural units, $R_{0}=R(0)$ and $T_{0}=R_{0}\sqrt{\rho/\Delta p}$, Equation (10) becomes Equation (9) with $\gamma=-4$. Thus Equation (10) is equivalent to Equation (1) with $\gamma=-4$ and $k=R_{0}^{3}\Delta p/\rho$.
Figure 1: Potential $\phi(r)$, given in Equation (12), for different values of $\gamma$, equally spaced by $\Delta\gamma=0.2$. There are qualitatively distinct families at $\gamma<-1$ (red–yellow) and $\gamma>-1$ (blue–green), as discussed in Section IV.1. They are separated by an isolated critical curve for $\gamma=-1$. Thin black lines highlight the coloured curves with a printed $\gamma$-value.
II.2.3 Lagrangian
@@
The Euler-Lagrange equation, $\tfrac{{\rm d}}{{\rm d}t}(\partial\mathcal{L}/\partial\dot{r})=\partial\mathcal{L}/\partial r$, then generates Equation (5), for all $\gamma$.
Selected potentials $\phi(r)$ are shown in Figure 1. The existence of two regimes, separated by the critical value $\gamma=-1$, is the fundamental reason for the qualitatively different behaviour of $r(t)$ in these regimes. These regimes will be discussed in more detail in Section IV.1.
III General solution
In this section, I present compact general solutions of Equation (1) with initial conditions of Equation (2), following an approach similar to [9], but for general real parameters $\gamma$. For convenience, all derivations and solutions are presented for the dimensionless form in Equation (5) with initial conditions stated in Equation (6). Any solution $r(t)$ can be transformed to the corresponding dimensional solution $R(T)$ via the linear transformations of Equation (3).
III.1 Inverse solution $t(r)$
III.1.1 Cases $\gamma\neq-1$
This case can be solved by reduction to quadrature, a standard method expressing the solution of ODEs in terms of integrals. For general second order ODEs, $\ddot{f}(t)=f(\dot{r},r,t)$, the integrals found in this way must be evaluated numerically. Interestingly, however, the limited form of Equation (5) allows us, through adequate substitutions, to express the emerging integrals through the incomplete beta function $B(x;\alpha,\beta)$ (derivation in Appendix A). This special function benefits from an extensive mathematical literature available for further analytical manipulations of the solution. Moreover, by introducing the shorthands
@@ III.1.1 Cases $\gamma\neq-1$
and
$\alpha=\frac{1}{4}+\frac{3-\gamma}{4|1+\gamma|},$ (14)
it turns out possible to compact the explicit solution $r(t)$ to a universal form, valid for all $\gamma\neq 1$, hence simplifying and generalising state-of-the-art hypergeometric solutions for $\gamma=-4$ [9] and lower negative integer $\gamma$ [10].
Explicitly, Equations (5) and (6) solve to
$t(r)=\tau-\sqrt{\frac{\eta}{2}}~{}B\left(r^{|1+\gamma|};\alpha,\tfrac{1}{2}\right),$ (15)
following the derivation provided in Appendix A. The collapse time $\tau$ is obtained by solving Equation (15) for $\tau$ at $(r,t)=(1,0)$,
$\tau(\gamma)=\sqrt{\frac{\eta}{2}}~{}B\big{(}\alpha,\tfrac{1}{2}\big{)}.$ (16)
Here $B(\alpha,\beta)$ is the (complete) beta function, i.e., the incomplete $B(x;\alpha,\beta)$ evaluated at $x=1$. Selected explicit values of $\tau$ are listed in Table 2.
Substituting Equation (16) back into Equation (15), the latter can be transformed to the special solution for $\gamma>-1$ found in modelling polytropic gas densities in galactic disks [13].
III.1.2 Special case $\gamma=-1$
In this case, Equation (8b) integrates to
$t(r)=\tau(-1)\,{\rm erf}\Big{(}\sqrt{-\ln r}\,\Big{)}$ (17)
where
$\tau(-1)=\sqrt{\pi/2}$ (18)
is the collapse time, reached for $r\rightarrow 0_{+}$. This collapse time is equal to the limit of Equation (16) for $\gamma\rightarrow-1$. Likewise, Equation (17) is the limit of Equation (15) for $\gamma\rightarrow-1$, showing that these singularities of Equations (15) and (16) at $\gamma=-1$ are removable.
III.2 Explicit solution $r(t)$
To invert the solutions $t(r)$ it is convenient to normalise Equation (15) by $\tau$, which yields
$\frac{t}{\tau}=1-I\left(r^{|1+\gamma|};\alpha,\tfrac{1}{2}\right),$ (19)
where $I(x;\alpha,\beta)\equiv B(x;\alpha,\beta)/B(\alpha,\beta)$ is called the regularised incomplete beta function (a simple, but useful step pointed out by Roberto Iacono, priv. com.). Equations (19) and (17) can be readily inverted to a compact general explicit solution in terms of well-known special functions,
$r(t)=$ $\exp\left(-{\rm erfi}^{2}\Big{(}\sqrt{2/\pi}~{}t\Big{)}\right)$ if $\gamma\!=\!-1$ , (20b)
where ${\rm erfi}(x)$ is the inverse error function and $Q(x;\alpha,\beta)$ is the inverse regularised incomplete beta function. Equation (20b) is the limit of Equation (20a) for $\gamma\rightarrow-1$, again showing that this singularity of Equation (20a) is of a removable type.
The absolute value $|t|$ in Equation (20a) does not follow from Equation (19), only valid for $t\in[0,\tau]$. It is an ad-hoc extension of $r(t)$ to the domain $t\in[-\tau,\tau]$, exploiting the time-reversal symmetry of Equations (5) and (6). Equation (20b) is already time-symmetric and hence valid on $t\in[-\tau,\tau]$.
The significance of Equation (20b) relies in the fact that the regularised incomplete beta function $I(x;\alpha,\beta)$ is the cumulative density function of the beta distribution, a common distribution function in probability theory, related to, but not to be confused with, the beta function. Hence, the function $Q(x;\alpha,\beta)$ is the quantile function of the beta distribution. This connection to the beta distribution, which has been intensively studied in probability theory, makes Equation (20b) a practical solution for further analytical manipulations. Moreover, the quantile function $Q(x;\alpha,\beta)$ is readily accessible to all modern programming languages, making this solution easy to implement. Two examples in Python and R are given in Appendix B.
Figure 2 shows $r(t)$ on the interval $t\in[0,\tau(\gamma)]$ for all integer $\gamma$ from $-4$ to $+4$. Key properties of these solutions will be discussed in Section IV.
@@ III.2 Explicit solution $r(t)$
Table 2: List of dimensionless collapse times $\tau$, given by Equations (16) and (18), and dimensionless collapse point velocities $\dot{r}(\tau)$, given by Equation (23). Exact solutions are given where they are algebraic, at least up to a factor $\pi$.